The population proportion is 0.50. What is the probability that a sample proportion will be within ±0.04 of the population proportion for each of the following sample sizes? Round your answers to 4 decimal places. Use z-table.

A.) n=100
B.) n= 200
C.) n=500
D.) n=1,000
Answer #1

We know that 0.25 PN(0.5, 0.51-0.5 N(0.5, N(0.5, 2

A)

Required probability = 0.04 100

P(-0.8 < Z < 0.8) P(Z < 0.8) _ P(Z <-0.8)

2P(Z < 0.8) _ 1-2(0.7881)-1-0.5762

B)

Required probability = 0.04 200

P(-1.13 < Z < 113) P(Z < 1.13) _ P(Z <-1.13)

2P(Z < 1.13)-1 = 2(0.8708)-1 = 0.74 16

C)

Required probability = 0.04 500

P(-1.79 < Z < 1.79) = P(Z < 1.79)-P(Z <-1.79)

2P(Z < 1.79)-1 = 2(0.9633)-1 = 0.9266

D)

Required probability = P(|p-P|<0.04)=P(|Z|<rac{0.04}{sqrt{rac{0.25}{1000}}})=P(|Z|<2.53)

=P(-2.53<Z<2.53)=P(Z<2.53)-P(Z<-2.53)

2PlZ < 2.53)-1 = 2(0.9943)-1 = 0.9886